OpenStudy (lukecrayonz):

Your friend runs up to you, scared that he is not ready for the upcoming quadratics test. To help him study, you will create four different quadratic functions. Then demonstrate to him how to rewrite each function as a group of factors, if possible. The function f(x) is a difference of squares. The function g(x) is a sum of squares. The function h(x) is a perfect square trinomial. The function j(x) can only have a GCF factored out of it.

OpenStudy (lukecrayonz):

I just don't know the sum of squares part lol

OpenStudy (lncognlto):

If the difference of squares equals, say, (x+3)(x-3), then I think I think the sum of squares = (x+3)(x+3).

OpenStudy (lncognlto):

Delete the one "I think" out of that sentence... XD

OpenStudy (lukecrayonz):

Weird. Never heard of that.

OpenStudy (lukecrayonz):

What is a trinomial that only can have a GCF factored out and nothing else?

OpenStudy (lukecrayonz):

Nevermind I got one

OpenStudy (lncognlto):

Ja, I've never heard of the some of squares before either... >.> Oh, ok. XD I was busy looking it up just to make sure, lol.

OpenStudy (lukecrayonz):

I literally just wrote an equation huge 400x^2+200000x-50000 lol GCF 400

OpenStudy (lukecrayonz):

Given the function k(x) = x^2, compare and contrast how the application of a constant, c, affects the graph. The application of the constant must be discussed in the following manners: k(x + c) k(x) + c k(cx) c • k(x)

OpenStudy (lncognlto):

LOL But yeah, that looks correct...

OpenStudy (lncognlto):

Well, I can say for certain sure that in k(x) + c, the value of c determines the y intercept, so it shifts the graph up or down the y-axis.

OpenStudy (lukecrayonz):

Hmm... I knew that but idk about the others.

OpenStudy (lncognlto):

As for the others, maybe this helps? If y = A (Bx + C ) + D C = vertical shrink, by a factor of "C" B = vertical shrink, by a factor of "B" D = vertical shift, D > 0, up, D < 0, down

OpenStudy (lncognlto):

I found this other question on OS that is the same: http://openstudy.com/study#/updates/525452b3e4b037f058866156 jdoe seems to have put it well. xD

OpenStudy (lukecrayonz):

It's 4:40 a.m and I still hate math

OpenStudy (lukecrayonz):

I never understood why I agree to do peoples homework for free lol

OpenStudy (lncognlto):

LOL At 4:40 a.m., I hate math myself. xD I remember sitting up trying to do trig at 2 a.m. and it was awful xD But I'm finished maths for school now... ^-^

OpenStudy (lukecrayonz):

Lucky you, I'm in calc 3 then I have to finish up statistics then I have to do my grad level math courses:/

OpenStudy (lncognlto):

e.e That sounds... nasty. XD I have to take a mathematics course next year in university as well. Some days I dread this, and some I actually look forward to it. o.o LOL

OpenStudy (lukecrayonz):

Never look forward to math. Ever. My past math test was so hard I literally screamed of happiness that I got a 60

OpenStudy (lukecrayonz):

So stupid. They made one of my engineering classes like crazy grading. If you had a 35% in the class, you got an A. That means you only have to know 35% of the information?!

OpenStudy (lncognlto):

I wrote my final maths exam like three weeks ago, and I was like, oh, I hope I passed >.> I still don't know. xD 35%?! Darn... If I got less than a 50%, I failed. I had to know every single little nugget of information hidden between the lines in my book, plus what they never told me, but still expected me to know.

OpenStudy (lncognlto):

As well as I lost marks if I didn't set out my work just right.

OpenStudy (lukecrayonz):

Haha it's only for engineering classes. If not, the pass rate I think he said would be 1 in 800

OpenStudy (lncognlto):

XD Well, I must be going as I have three more upcoming exams to study for... e.e Hopefully that other question helped a bit... xD